CAVITY RINGDOWN SPECTROSCOPY WITH A CW CO_{2}-LASER,MICROWAVE-MODULATED SIDE-BAND SYSTEM

\underline{C.R. BUCHER} AND K.K. LEHMANN, Department of Chemistry, Princeton University, Princeton, NJ 08544; G.T. FRASER AND D.F. PLUSQUELLIC, Optical Technology Division, National Institute of Standards and Technology, Gaithersburg, MD 20899.

This work reports the first observation of a ringdown decay from an empty resonator using a < 0.5 MHz linewidth, cw, carbon dioxide laser with ∼ 1 mW of microwave sidebands separated 8 GHz to 18 GHz from the CO_{2} carrier frequency. The resonator consists of a Fabry-Perot interferometer with two highly reflective mirrors (R∼99.5%) separated by 1.2 m providing an optical pathlength for absorption of 420 m and a theoretical ringdown time of 800 ns. At present, the lack of availability of highly reflective mid-infrared mirrors places an upper limit on the ringdown time. One of the benefits of using an extremely high resolution laser is that a single mode of the cavity can be selectively monitored. Future investigations with the experimental system include saturation spectroscopy of unstable chemical species, such as ozone, for accurate concentration determination.